Two charges q1= 1.0 Coul and q2= 2.0 Coul are separated by



10.0 cm. At what point line joining the two charges is the electric field strength



is zero?

1
Expert's answer
2022-01-12T08:28:58-0500

Assume the field is zero at distance x from the second charge. The field is zero at he point where the magnitudes of the fields created by each individual charge are the same.

To find the point where the field is zero, write the appropriate equation:


"E_1=E_2,\\\space\\
\dfrac{1}{4\pi\epsilon_0}\dfrac {q_1}{(r-x)^2}=\dfrac{1}{4\pi\epsilon_0}\dfrac {q_2}{(x)^2},\\\space\\
\dfrac {q_1}{(r-x)^2}-\dfrac {q_2}{x^2}=0,\\\space\\
x_1=5.86\text{ cm}."